OpenStudy (anonymous):

Line a has an equation y=(m-3)x+3 , where m is any integer. Line b has an equation y=-x+4 What is the value of m if line a is parallel to line b? What is the value of m if line a is perpendicular to line b? For what values of m will the slope of line lie between 0 and 4 (not including either value)? Drop the numbers in increasing order (((the choices are 0-9)))

OpenStudy (mathmate):

L1: y=(m-3)x+3 L2: y=-x+4 @guppy83711 Can you tell me the slopes of L1 and L2?

OpenStudy (anonymous):

it dosnt say... it just says the slope is any integer @mathmate

OpenStudy (mathmate):

@guppy83711 I expect you to give me the slope of each of the lines. It is necessary to know these quantities to solve the questions.

OpenStudy (mathmate):

The slope of an equation of a line in slope-intercept form (same form as these equations) y=ax+b, then a is the slope, and b is the intercept. Now can you give me the slopes of L1 and L2?

OpenStudy (anonymous):

i dont get how to find line 1's slope but L2 is -1 @mathmate

OpenStudy (mathmate):

For L2, slope is -1. What you need to do is to look for the coefficient of x. If the coefficient is not a pure number, is is still the slope. For L1, the coefficient of x in L1: (m-3)x+3 is what's left when the term (m-3)x is stripped off the x, what's left is (m-3), and so (m-3) is the slope of L2. Remember we are working with algebra. Not all quantities are pure numbers. (m-3) is still a quantity which equals the slope. Is that more clear?

OpenStudy (anonymous):

Kind of... If (m-3) is the slope then how would i solve this?

OpenStudy (mathmate):

"What is the value of m if line a is parallel to line b?" If L1 and L2 are parallel, then their slopes are equal. Can you solve for m then?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

If they are parallel, the slopes are \(equal\). So -1=m-3, can you solve for m?

OpenStudy (anonymous):

m=2? @mathmate

OpenStudy (mathmate):

That is correct! Are you ready for the next question?

OpenStudy (anonymous):

yea @mathmate

OpenStudy (mathmate):

"What is the value of m if line a is perpendicular to line b?" When two lines are perpendicular to each other, the product (times) of the two slopes equals -1. i.e. (-1)(m-3)=-1 Can you solve for m ?

OpenStudy (anonymous):

-4?

OpenStudy (mathmate):

Hmmm let me see: (-1)(-4-3)=(-1)(-7)=7 \(\ne\) -1 so m=-4 is not correct. Can you try again, or show your work?

OpenStudy (anonymous):

(-1)(4-3)=(-1)(1)=-1? m=4?

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Yay, you got it again! Are you ready for the next one?

OpenStudy (anonymous):

Yes

OpenStudy (mathmate):

"For what values of m will the slope of line lie between 0 and 4 (not including either value)? Drop the numbers in increasing order (((the choices are 0-9)))" Now you need to solve for the slope of L1 two times: once slope=0, find m, and the other slope = 4, and find m. Then you would write them in an interval, like (2,6) with the numbers that you find. Get set, go!!

OpenStudy (anonymous):

What?? i just got really confused...

OpenStudy (mathmate):

Solve one at a time. If L1 has a slope of 0, what's m?

OpenStudy (anonymous):

I thought m was slope? @mathmate

OpenStudy (mathmate):

m is not slope of L1. (m-3) is slope of L1, and that equals 0. What is m?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Correct! Now go for slope of L1=4. What would m be?

OpenStudy (anonymous):

1? @mathmate

OpenStudy (mathmate):

(1-3)\(\ne\)4, so you'd have to try again! Sorry

OpenStudy (anonymous):

Oh it is 7. I thought it was addition for some reason @mathmate

OpenStudy (mathmate):

ok. So you got 1 and 7. We need to put the numbers in an interval, in the form (a,b), where a is the smaller number, and b is the larger number. We also have to exclude a and b (we were told to), so we put (a,b). If we want to include a, we would write [a,b), or both, then [a,b]. Use 1 and 7 to write your interval and you're done! :)